OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 3i

OpenStudy (jdoe0001):

recall that \(\bf \large (a-b)(a+b) = a^2-b^2\)

OpenStudy (anonymous):

for making o conjugate just convert i to -i so: 1+3i --> 1-3i (1+3i)(1-3i)=..

OpenStudy (haseeb96):

its conjugate is 1 - 3i and the product of it is (1=3i)(1-3i)=1-(3i)^2

OpenStudy (anonymous):

1 + 9i

OpenStudy (anonymous):

(1-3i)(1+3i)=1+3i-3i-9(i^2)=1+9=10 because i^2=-1

OpenStudy (haseeb96):

1-9(-1) since i^2=-1 1+9=10 that is